# A Two-Timescale Primal-Dual Framework for Reinforcement Learning via Online Dual Variable Guidance

### v2

@@ 3 Synchronous Algorithm
We now propose our solution to the saddle-point problem (9) that is based on a projected stochastic gradient descent-ascent method summarized in Algorithm 1. The algorithm does not assume to know the transition probabilities of the MDP but has access to a generative model [Kakade, 2003, Chen and Wang, 2016], also known as a sampling oracle, which provides sampling access to any state in the environment. Given a state-action pair $(s,a)$ as input, the generative model provides the reward $r(s,a)$ and a follow-up state $s^{\prime}$ with $s^{\prime}\sim\mathcal{P}(\cdot|s,a)$. The almost sure convergence analysis is based on the two-timescale stochastic approximation framework developed in Borkar [1997], Borkar [2023]. The dynamics of the projected recursion are based on results of Dupuis [1987] and Dupuis and Nagurney [1993].
Since the true gradients of the Lagrangian $L$ depend on the unknown transition probabilities, we use stochastic gradient approximations, denoted by $\widehat{\nabla}_{V}L$ and $\widehat{\nabla}_{\rho}L$, to perform the updates. To ensure that the iterates remain bounded, we project the dual iterates onto the feasible set $H$ using the bounds on the optimal dual variable established in Lemma 2.1. In the synchronous setting, we fully update the primal and dual variables in each iteration. We construct the stochastic gradients from generative model transition samples obtained for each state-action pair.
@@
Algorithm 1 is a stochastic approximation algorithm based on the primal-dual gradient method for saddle-point estimation [Arrow et al., 1958]. Chen and Wang [2016], Lee and He [2019], among others, study versions of a related approach for unregularized LP formulations, where the optimal policy’s uniqueness must be assumed. They provide sample complexity bounds by applying a concentration bound to the average iterate. In contrast, we take the dynamical systems viewpoint [Borkar, 2023] on the stochastic approximation algorithm and analyze the asymptotic almost sure convergence. This change in perspective is reflected in the choice of stepsizes, where our approaches feature diminishing stepsizes operating on two timescales, such that the last iterates asymptotically approach the limiting ODEs. In contrast, they rely on a fixed stepsize scheme, which suffices for the average iterate analysis.
@@ 3.1 Almost Sure Convergence Analysis of Algorithm 1
Theorem 3.1 (Convergence of Algorithm 1).
@@
Examples of suitable stepsize sequences are $\alpha_{k}=\frac{1}{k^{q}},q\in(\frac{1}{2},1)$ and $\beta_{k}=\frac{1}{k}$ or $\alpha_{k}=\frac{1}{k}$, $\beta_{k}=\frac{1}{1+k\log k}$. The proof of Theorem 3.1 builds on the two-timescale stochastic approximation framework of Borkar [1997]. This framework takes the dynamical systems viewpoint on the stochastic iteration and applies to the regularized MDP problem addressed with Algorithm 1 as follows.
@@
Proposition 3.1 (Gradient and noise properties of Algorithm 1).
The following conditions are satisfied.
@@
Together with the two-timescale stepsize sequences assumed in Theorem 3.1, these gradient and noise properties ensure that the coupled system (12) asymptotically approximates a dynamical system with a fast transient $V(t)$ and a slowly evolving dual variable $\rho(t)$. The separation of the timescales allows the fast transient to view the dual variable as quasi-constant, and in turn, the dual variable to assume the transient to be equilibrated. The convergence of the stochastic recursion is derived from the asymptotic properties of the limiting dynamical system. The ODE describing the dynamics of the primal variable $V$ is the gradient flow
@@
The dual variable’s dynamic is restricted to the compact set $H$. We follow the notation of [Kushner and Yin, 2003, Section 4.3] for the projected dynamics. For $\rho\in\partial H$, the boundary of $H$, let $C(\rho)$ be the infinite convex cone generated by the outer normals at $\rho$ of the faces on which $\rho$ lies. For $\rho$ in the interior of $H$, let $C(\rho)=\{0\}$. The projected ODE describing the flow of the dual variable is then given by
@@
Proposition 3.2 (Dynamical systems properties of Algorithm 1).
The following conditions are satisfied.
@@
    Let $\{(V_{k},\rho_{k})\}$ be a sequence of iterates generated by (12) with a stepsize sequence fulfilling the assumption of Theorem 3.1. Then $\sup_{k}(\|V_{k}\|+\|\rho_{k}\|)<\infty,$ a.s.
@@
Given the above stability properties, existing stochastic approximation theory by Borkar [2023] ensures the asymptotic convergence as follows.
Proof of Theorem 3.1
Given Propositions 3.1, and 3.2, and the two-timescale stepsize sequences assumed in Theorem 3.1, [Borkar, 2023, Theorem 8.1] applies and the coupled iterates $(V_{k},\rho_{k})$ generated by the recursions (12), and hence by Algorithm 1, converge -$\mathbb{P}\text{-a.s.}$ to $(\lambda(\rho^{\star}),\rho^{\star})$. $\blacksquare$
@@ 4 Asynchronous Algorithm
Assuming access to a generative model, as in Algorithm 1, can be restrictive in applications where samples can only be obtained by interacting with the environment. We therefore extend the approach to the standard online setting in which data are generated along a trajectory induced by a (fixed or time-varying) behavioral policy. This covers classical online methods such as Q-learning [Watkins and Dayan, 1992]. The fixed behavioral policy setting is also related to offline RL where policies are learned from a fixed set of past experiences, see Levine et al. [2020] for an overview and Gabbianelli et al. [2024], Hong and Tewari [2024] for LP-based formulations.
Our algorithm retains the projected stochastic gradient descent-ascent structure, but performs asynchronous updates: only the state-action components visited by the behavioral policy are updated at each iteration. The almost-sure convergence analysis follows the asynchronous stochastic approximation framework of Konda and Borkar [1999], Perkins and Leslie [2013], while our finite-time rate bounds build on the two-timescale analysis developed in Zeng et al. [2024].
In addition to the related works mentioned in the discussion of Algorithm 1, the asynchronous algorithm shares features with the algorithms SBEED [Dai et al., 2018b] and Dual-AC [Dai et al., 2018a]. They approximate different saddle-point formulations of the RL problem with behavioral policy access to the stochastic environment by sampling transitions from a growing experience replay buffer. In their convergence analysis, they assume a fixed behavioral policy. Our proposed algorithm combines the experience replay-based gradient estimation with the separation of the nested loop minimization by the two-timescale method and on-policy exploration. This combination allows us to apply the asynchronous stochastic approximation analysis in a way that has not been proposed for the regularized LP approach to reinforcement learning.
@@
At each iteration, we add the current transition of the behavioral policy to the experience replay buffer, denoted as $\mathcal{D}_{k}$ in iteration $k$. To efficiently leverage the gathered experience, we opt for a structured buffer with a separate list $\mathcal{D}_{k}(s,a)$ for each state-action pair $(s,a)$. Furthermore, for each state $s^{\prime}$, we maintain a set [^1] The set of incoming state-action pairs does not contain duplicates. In contrast, the replay buffer consists of lists that allow duplicate entries., denoted as $\mathcal{D}_{\mathrm{inc}}(s^{\prime})$, containing the “incoming" state-action pairs that have previously led to state $s^{\prime}$. Hence, if $X_{k}\notin\mathcal{D}_{\mathrm{inc}}(s_{k})$, we update $\mathcal{D}_{\mathrm{inc}}(s_{k}):=\mathcal{D}_{\mathrm{inc}}(s_{k})\cup\{X_{k}\}$.
@@ 4.1 Almost Sure Convergence Analysis of the Algorithm 2
Assumption 2 (Asynchronous Step-Size Sequences).
@@
Theorem 4.1 (Almost Sure Convergence of Algorithm 2).
Consider the stochastic gradients defined in (14). Let Assumption 1 hold and choose stepsize sequences $\{\alpha_{k}\}$$\{\beta_{k}\}$ that satisfy Assumption 2. Let $\{(V_{k},\rho_{k})\}$ be a sequence generated by Algorithm 2 either with on-policy exploration or with a fixed strictly exploratory behavioral policy. Then, the last iterates converge $\mathbb{P}$-a.s. to the regularized saddle-point of (9), that is,
@@
The proof of Theorem 4.1 is based on the asynchronous stochastic approximation theory by Borkar [1998], Konda and Borkar [1999], and the generalization in Perkins and Leslie [2013] that clarifies the conditions on the update frequencies.
@@
Similarly to the synchronous case, the asymptotic analysis of the asynchronous stochastic iteration consists of analyzing the equilibrium properties of the dynamical systems that describe the limiting behavior of the iteration. In the asynchronous setting, the random selection of update components impact the limiting dynamics. Following the notation of Perkins and Leslie [2013], define for general $\epsilon>0,d\in\mathbb{N}$, $\Omega^{\epsilon}_{d}:=\{\text{diag}(w_{1},\dots,w_{d})|w_{i}\in[\epsilon,1],i=1,\dots,d\}$. For the fast iterate $V$, we are concerned with the ODE
@@
The properties (B1)-a,b and (B4) do not change compared to the respective assumptions for the synchronous scheme. Under property (B1)-c, the bias term $\mathcal{E}_{k}$ vanishes and does not affect the limiting dynamics (see Remark 3 in [Borkar, 2023, Section 2.2]). Properties (B2) and (B4) are the asynchronous counterparts of (A4) and (A5. Property (B5) ensures that the state-action index process $(s_{k},a_{k})$ is an ergodic Markov chain on $\mathcal{X}$, which implies that every state-action component of $\rho$ and every state component of $V$ is updated with positive asymptotic frequency (see also [Perkins and Leslie, 2013, Lemma 3.4]).
Proof of Theorem 4.1
Let Assumptions 1 and 2 hold, and let the stochastic gradients be as stated in (14). Given the properties listed in Proposition 4.1 and with the asynchronous step sizes defined in Assumption 2, the fast iterates’ interpolated trajectory converges to the trajectory of a solution of the differential inclusion (19a) [Perkins and Leslie, 2013, Lemma 4.3], and the slow iterate asymptotically tracks a solution to (19b) almost surely [Perkins and Leslie, 2013, Theorem 4.7].
Finally, the almost sure convergence of the last iterates of Algorithm 2 follows from the global asymptotic equilibrium property of the projected ODE (18b), (B3), by applying [Perkins and Leslie, 2013, Corollary 4.8]. $\blacksquare$
@@ 4.2 Convergence Rate Analysis of Algorithm 2
As discussed in the proof of Corollary 2.1, this projection does not affect the optimal primal variable.
To establish the almost sure convergence result in Theorem 4.1, we assumed that for any strictly exploratory policy, the induced Markov chain is ergodic (Assumption 1). For the derivation of the convergence rate, we strengthen this assumption slightly as follows.
Assumption 3 (Deterministic uniformly geometrically ergodic policy).
There exists a deterministic state-feedback policy $\pi_{d}:\mathcal{S}\to\mathcal{A}$ such that the Markov chain with kernel $\mathcal{P}_{d}(\cdot|s):=\mathcal{P}(\cdot|s,\pi_{d}(s))$ is uniformly geometrically ergodic.
Because the policies $\pi_{\rho}$ induced by the dual variables are strictly exploratory with a uniform lower bound $\pi_{\rho}(a|s)\geq\tfrac{C^{L}}{|\mathcal{A}|C^{U}}>0$, this strengthened assumption ensures the uniform geometric ergodicity for all dual-induced chains. We denote the one-step kernel under a fixed parameter $\rho\in H$ as
$\mathcal{P}_{\rho}\big((s^{\prime},a^{\prime})|(s,a)\big):=\pi_{\rho}(a^{\prime}|s^{\prime})\mathcal{P}(s^{\prime}|s,a),$
with the dual-induced policy $\pi_{\rho}$, defined in (3).
Lemma 4.1 (UGE of dual policy chains).
Let Assumption 3 hold. Then, the family of Markov chains corresponding to the transition kernels $\{\mathcal{P}_{\rho}\}_{\rho\in H}$ is uniformly geometrically ergodic, with common constants for all $\rho\in H$. That is, there exist common constants $C_{\mathcal{X}}>0$ and $\varrho\in(0,1)$, such that, for all $\rho\in H$, the Markov chain with kernel $\mathcal{P}_{\rho}$ has a unique stationary law $\mu_{\rho}$ and for any $k\geq 0$,
$\sup_{x\in\mathcal{X}}\|\delta_{x}\mathcal{P}_{\rho}^{k}-\mu_{\rho}\|_{\mathrm{TV}}\leq C_{\mathcal{X}}\varrho^{k}.$
Note that $\mu_{\rho}(s,a)=\tilde{\mu}_{\rho}(s)\pi_{\rho}(a|s)$ holds, with $\tilde{\mu}_{\rho}(s)=\sum_{a\in\mathcal{A}}\mu_{\rho}(s,a)$ the state marginal of $\mu_{\rho}$. The proof, stated in Appendix C.1, is based on the observation that for any $\rho\in H$ there is a uniformly positive probability of following the ergodic deterministic policy, due to the strictly positive uniform lower bound on the action probabilities $\pi_{\rho}(a|s)$ across states and actions.
The convergence rate analysis presented here builds on the recent work of Zeng et al. [2024]. The analysis applies with a specific choice of the two-timescale stepsize sequences. In the following let the stepsize sequences be
@@
We extend and adapt the convergence rate analysis of Zeng et al. [2024] for our LP-based algorithm framework and asynchronous updating scheme in two important nontrivial aspects.
First, while Zeng et al. [2024] considers only the global decaying stepsize setting, our focus is on the asynchronous regime with local stepsizes, which introduces randomness and non-monotonicity to the realized stepsize sequence. To handle the non-monotonicity, we switch the analysis to the currently least frequently visited state or state-action. This coordinate still receives updates linear in iteration counter $k$ eventually, with a high-probability bound on the time when the linear growth rate applies. The proof is based on the uniform geometric ergodicity of the dual policy induced chains and Lipschitz properties of the dual-induced Markov kernels.
Second, unlike Zeng et al. [2024], the stochastic gradient estimator based on the replay buffer is biased in our setting. We find that the bias decays at a rate fast enough to not become dominating.
These properties of the asynchronous PGDA iterates as well as further regularity and boundedness properties form the basis of our convergence rate proof and are stated in the following proposition. Throughout, let $p_{\star}$ denote a uniform lower bound on the dual-variable induced stationary distributions over the state-action space, valid for all $\rho\in H$. [^3] An explicit expression for $p_{\star}$ is derived in the proof of Lemma C.1. Denote the minimal number of visits to any state-action pair until iteration k as $\underline{\nu}_{k}:=\min_{(s,a)\in\mathcal{X}}\nu_{k}(s,a)$. It holds that $\underline{\nu}_{k}\leq\nu_{k}(s,a)\leq k$ for all $(s,a)\in\mathcal{X}$. Denote the reduced Lagrangian objective as
@@
Recall that $\bar{\alpha}_{k}=\alpha(\tilde{\nu}_{k}(s_{k}))$ and $\bar{\beta}_{k}=\beta(\nu(X_{k}))$.
Proposition 4.2.
@@
The proof of Proposition 4.2 is provided in Section C.1.
@@
$\tau_{k}^{2}\alpha_{\mathrm{env},k-\tau_{k}}\leq\tfrac{\tilde{\eta}_{V}}{18\,C_{1}\,C_{\mathrm{env}}},\quad\tau_{k}\alpha_{\mathrm{env},k-\tau_{k}}\leq\min\{\tfrac{1}{\sqrt{3}D},\,\tfrac{1}{\sqrt{3}LB}\},\quad\alpha_{k}\leq\tfrac{1}{\tilde{\eta}_{V}},\quad\tfrac{p_{\star}k}{2}\geq 1,$ (24)
@@
Theorem 4.2 (Convergence rate of Algorithm 2).
Let Assumption 3 hold. Let $\{\rho_{k}\}$, $\{V_{k}\}$ denote the iterates of Algorithm 2 with the projection on the primal iterates to $\mathcal{V}_{r}$. Fix $\delta\in(0,1)$ and let $\mathcal{G}_{\delta}$, and $K(\delta)$ be as in Proposition 4.2. Let the stepsize schedules be as in (21) with
$\beta_{0}\geq\frac{2}{\mu_{\mathrm{opt}}},\ \ \frac{\beta_{0}}{\alpha_{0}}\leq\frac{\tilde{\eta}_{V}}{8\mu_{\mathrm{opt}}}.$
@@
$\mathbb{E}\left[f(\rho^{\star})-f(\rho_{k})|\mathcal{G}_{\delta}\right]$ $\leq\left(\mathbb{E}\left[f(\rho^{\star})-f(\rho_{k_{0}})|\mathcal{G}_{\delta}\right]+\frac{16L^{2}}{p_{\star}\tilde{\eta}_{V}}\frac{\beta_{k_{0}}}{\alpha_{k_{0}}}\mathbb{E}\left[\|V_{k_{0}}-\lambda(\rho_{k_{0}})\|_{2}^{2}|\mathcal{G}_{\delta}\right]\right)\frac{k_{0}+1}{k+1}$
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ +\frac{2C_{2}(\delta)\log^{2}(k+1)}{3(k+1)^{2/3}},$

### v3

@@ 3 Synchronous Algorithm
We now propose our solution to the saddle-point problem (9) that is based on a projected stochastic gradient descent-ascent method summarized in Algorithm 1. The algorithm does not assume to know the transition probabilities of the MDP but has access to a generative model (Kakade, 2003; Chen and Wang, 2016), also known as a sampling oracle, which provides sampling access to any state in the environment. Given a state-action pair $(s,a)$ as input, the generative model provides the reward $r(s,a)$ and a follow-up state $s^{\prime}$ with $s^{\prime}\sim\mathcal{P}(\cdot|s,a)$. The almost sure convergence analysis is based on the two-timescale stochastic approximation framework developed in Borkar (1997), Borkar (2023). The dynamics of the projected recursion are based on results of Dupuis (1987) and Dupuis and Nagurney (1993).
Since the true gradients of the Lagrangian $L$ depend on the unknown transition probabilities, we use stochastic gradient approximations, denoted by $\widehat{\nabla}_{V}L$ and $\widehat{\nabla}_{\rho}L$, to perform the updates. To ensure that the iterates remain bounded, we project the dual iterates onto the feasible set $H$ using the bounds on the optimal dual variable established in Lemma 5. In the synchronous setting, we fully update the primal and dual variables in each iteration. We construct the stochastic gradients from generative model transition samples obtained for each state-action pair.
@@
Algorithm 1 is a stochastic approximation algorithm based on the primal-dual gradient method for saddle-point estimation (Arrow et al., 1958). Chen and Wang (2016), Lee and He (2019), among others, study versions of a related approach for unregularized LP formulations, where the optimal policy’s uniqueness must be assumed. They provide sample complexity bounds by applying a concentration bound to the average iterate. In contrast, we take the dynamical systems viewpoint (Borkar, 2023) on the stochastic approximation algorithm and analyze the asymptotic almost sure convergence. This change in perspective is reflected in the choice of stepsizes, where our approaches feature diminishing stepsizes operating on two timescales, such that the last iterates asymptotically approach the limiting ODEs. In contrast, they rely on a fixed stepsize scheme, which suffices for the average iterate analysis.
@@ 3.1 Almost Sure Convergence Analysis of Algorithm 1
Theorem 6 (Convergence of Algorithm 1)
@@
Examples of suitable stepsize sequences are $\alpha_{k}=\frac{1}{k^{q}},q\in(\frac{1}{2},1)$ and $\beta_{k}=\frac{1}{k}$ or $\alpha_{k}=\frac{1}{k}$, $\beta_{k}=\frac{1}{1+k\log k}$. The proof of Theorem 6 builds on the two-timescale stochastic approximation framework of Borkar (1997). This framework takes the dynamical systems viewpoint on the stochastic iteration and applies to the regularized MDP problem addressed with Algorithm 1 as follows.
@@
Proposition 7 (Gradient and noise properties of Algorithm 1)
The following
conditions are satisfied.
@@
Together with the two-timescale stepsize sequences assumed in Theorem 6, these gradient and noise properties ensure that the coupled system (12) asymptotically approximates a dynamical system with a fast transient $V(t)$ and a slowly evolving dual variable $\rho(t)$. The separation of the timescales allows the fast transient to view the dual variable as quasi-constant, and in turn, the dual variable to assume the transient to be equilibrated. The convergence of the stochastic recursion is derived from the asymptotic properties of the limiting dynamical system. The ODE describing the dynamics of the primal variable $V$ is the gradient flow
@@
The dual variable’s dynamic is restricted to the compact set $H$. We follow the notation of (Kushner and Yin, 2003, Section 4.3) for the projected dynamics. For $\rho\in\partial H$, the boundary of $H$, let $C(\rho)$ be the infinite convex cone generated by the outer normals at $\rho$ of the faces on which $\rho$ lies. For $\rho$ in the interior of $H$, let $C(\rho)=\{0\}$. The projected ODE describing the flow of the dual variable is then given by
@@
Proposition 8 (Dynamical systems properties of Algorithm 1)
The following
conditions are satisfied.
@@
    Let $\{(V_{k},\rho_{k})\}$ be a sequence of iterates generated by (12) with a stepsize sequence fulfilling the assumption of Theorem 6. Then $\sup_{k}(\|V_{k}\|+\|\rho_{k}\|)<\infty,$ a.s.
@@
Given the above stability properties, existing stochastic approximation theory by Borkar (2023) ensures the asymptotic convergence as follows.
Proof of Theorem 6
Given Propositions 7, and 8, and the two-timescale stepsize sequences assumed in Theorem 6, (Borkar, 2023, Theorem 8.1) applies and the coupled iterates $(V_{k},\rho_{k})$ generated by the recursions (12), and hence by Algorithm 1, converge -$\mathbb{P}\text{-a.s.}$ to $(\lambda(\rho^{\star}),\rho^{\star})$. $\blacksquare$
@@ 4 Asynchronous Algorithm
Assuming access to a generative model, as in Algorithm 1, can be restrictive in applications where samples can only be obtained by interacting with the environment. We therefore extend the approach to the standard online setting in which data are generated along a trajectory induced by a (fixed or time-varying) behavioral policy. This covers classical online methods such as Q-learning (Watkins and Dayan, 1992). The fixed behavioral policy setting is also related to offline RL where policies are learned from a fixed set of past experiences, see Levine et al. (2020) for an overview and Gabbianelli et al. (2024); Hong and Tewari (2024) for LP-based formulations.
Our algorithm retains the projected stochastic gradient descent-ascent structure, but performs asynchronous updates: only the state-action components visited by the behavioral policy are updated at each iteration. The almost-sure convergence analysis follows the asynchronous stochastic approximation framework of Konda and Borkar (1999); Perkins and Leslie (2013), while our finite-time rate bounds build on the two-timescale analysis developed in Zeng et al. (2024).
In addition to the related works mentioned in the discussion of Algorithm 1, the asynchronous algorithm shares features with the algorithms SBEED (Dai et al., 2018b) and Dual-AC (Dai et al., 2018a). They approximate different saddle-point formulations of the RL problem with behavioral policy access to the stochastic environment by sampling transitions from a growing experience replay buffer. In their convergence analysis, they assume a fixed behavioral policy. Our proposed algorithm combines the experience replay-based gradient estimation with the separation of the nested loop minimization by the two-timescale method and on-policy exploration. This combination allows us to apply the asynchronous stochastic approximation analysis in a way that has not been proposed for the regularized LP approach to reinforcement learning.
@@
At each iteration, we add the current transition of the behavioral policy to the experience replay buffer, denoted as $\mathcal{D}_{k}$ in iteration $k$. To efficiently leverage the gathered experience, we opt for a structured buffer with a separate list $\mathcal{D}_{k}(s,a)$ for each state-action pair $(s,a)$. Furthermore, for each state $s^{\prime}$, we maintain a set, denoted as $\mathcal{D}_{\mathrm{inc}}(s^{\prime})$, containing the “incoming” state-action pairs that have previously led to state $s^{\prime}$. [^1] The set of incoming state-action pairs does not contain duplicates. In contrast, the replay buffer consists of lists that allow duplicate entries. Hence, if $X_{k}\notin\mathcal{D}_{\mathrm{inc}}(s_{k})$, we update $\mathcal{D}_{\mathrm{inc}}(s_{k}):=\mathcal{D}_{\mathrm{inc}}(s_{k})\cup\{X_{k}\}$.
@@ 4.1 Almost Sure Convergence Analysis of the Algorithm 2
Assumption 10 (Asynchronous Step-Size Sequences)
@@
Theorem 11 (Almost Sure Convergence of Algorithm 2)
Consider the stochastic
gradients defined in (14). Let Assumption 9 hold and choose stepsize sequences $\{\alpha_{k}\}$$\{\beta_{k}\}$ that satisfy Assumption 10. Let $\{(V_{k},\rho_{k})\}$ be a sequence generated by Algorithm 2 either with on-policy exploration or with a fixed strictly exploratory behavioral policy. Then, the last iterates converge $\mathbb{P}$-a.s. to the regularized saddle-point of (9), that is,
@@
The proof of Theorem 11 is based on the asynchronous stochastic approximation theory by Borkar (1998), Konda and Borkar (1999), and the generalization in Perkins and Leslie (2013) that clarifies the conditions on the update frequencies.
@@
Similarly to the synchronous case, the asymptotic analysis of the asynchronous stochastic iteration consists of analyzing the equilibrium properties of the dynamical systems that describe the limiting behavior of the iteration. In the asynchronous setting, the random selection of update components impact the limiting dynamics. Following the notation of Perkins and Leslie (2013), define for general $\epsilon>0,d\in\mathbb{N}$, $\Omega^{\epsilon}_{d}:=\{\text{diag}(w_{1},\dots,w_{d})|w_{i}\in[\epsilon,1],i=1,\dots,d\}$. For the fast iterate $V$, we are concerned with the ODE
@@
The properties (B1)-a,b and (B4) do not change compared to the respective assumptions for the synchronous scheme. Under property (B1)-c, the bias term $\mathcal{E}_{k}$ vanishes and does not affect the limiting dynamics (see Remark 3 in (Borkar, 2023, Section 2.2)). Properties (B2) and (B4) are the asynchronous counterparts of (A4) and (A5. Property (B5) ensures that the state-action index process $(s_{k},a_{k})$ is an ergodic Markov chain on $\mathcal{X}$, which implies that every state-action component of $\rho$ and every state component of $V$ is updated with positive asymptotic frequency (see also (Perkins and Leslie, 2013, Lemma 3.4)).
Proof of Theorem 11
Let Assumptions 9 and 10 hold, and let the stochastic gradients be as stated in (14). Given the properties listed in Proposition 12 and with the asynchronous step sizes defined in Assumption 10, the fast iterates’ interpolated trajectory converges to the trajectory of a solution of the differential inclusion (19a(Perkins and Leslie, 2013, Lemma 4.3), and the slow iterate asymptotically tracks a solution to (19b) almost surely (Perkins and Leslie, 2013, Theorem 4.7).
Finally, the almost sure convergence of the last iterates of Algorithm 2 follows from the global asymptotic equilibrium property of the projected ODE (18b), (B3), by applying (Perkins and Leslie, 2013, Corollary 4.8). $\blacksquare$
@@ 4.2 Convergence Rate Analysis of Algorithm 2
As discussed in the proof of Corollary 2, this projection does not affect the optimal primal variable.
To establish the almost sure convergence result in Theorem 11, we assumed that for any strictly exploratory policy, the induced Markov chain is ergodic (Assumption 9). For the derivation of the convergence rate, we strengthen this assumption slightly as follows.
We denote Dobrushin’s ergodic coefficient for a Markov kernel $Q$ on $\mathcal{X}$ as
It measures the maximal contraction of the kernel in total variation distance; see Dobrushin (1956) and Brémaud (2020, Definition 4.3.11).
Assumption 13 (Uniform mixing of dual-induced chains)
Let $\{\mathcal{P}_{\rho}\}_{\rho\in H}$ denote the
family of state-action transition kernels induced by the dual-induced policies $\{\pi_{\rho}\}_{\rho\in H}$.
  1. (i)
    There exist constants $C_{\mathcal{X}}>0$ and $\varrho\in(0,1)$ such that, for every $\rho\in H$, the Markov chain with kernel $\mathcal{P}_{\rho}$ has a unique stationary law $\mu_{\rho}$ and
    $\sup_{x\in\mathcal{X}}\|\delta_{x}\mathcal{P}_{\rho}^{k}-\mu_{\rho}\|_{\mathrm{TV}}\leq C_{\mathcal{X}}\varrho^{k},\qquad k\geq 0.$
  2. (ii)
    There exist constants $m_{\star}\in\mathbb{N}$ and $\kappa\in(0,1)$ such that, for every sequence $\{\rho_{t}\}_{t\in\mathbb{N}}\subset H$,
Remark 14
A sufficient condition for Assumption 13 is the existence of a deterministic state-feedback policy $\pi_{d}:\mathcal{S}\rightarrow\mathcal{A}$ whose induced state chain is uniformly geometrically ergodic, that is, it has a unique stationary law and converges to it geometrically fast in total variation, uniformly over initial states; see Lemma 24.
The convergence rate analysis presented here builds on the recent work of Zeng et al. (2024). The analysis applies with a specific choice of the two-timescale stepsize sequences. In the following let the stepsize sequences be
@@
We extend and adapt the convergence rate analysis of Zeng et al. (2024) for our LP-based algorithm framework and asynchronous updating scheme in two important nontrivial aspects.
First, while Zeng et al. (2024) considers only the global decaying stepsize setting, our focus is on the asynchronous regime with local stepsizes, which introduces randomness and non-monotonicity to the realized stepsize sequence. To handle the non-monotonicity, we switch the analysis to the currently least frequently visited state or state-action. This coordinate still receives updates linear in iteration counter $k$ eventually, with a high-probability bound on the time when the linear growth rate applies. The proof is based on the uniform geometric ergodicity of the dual policy induced chains and Lipschitz properties of the dual-induced Markov kernels.
Second, unlike Zeng et al. (2024), the stochastic gradient estimator based on the replay buffer is biased in our setting. We find that the bias decays at a rate fast enough to not become dominating.
These properties of the asynchronous PGDA iterates as well as further regularity and boundedness properties form the basis of our convergence rate proof and are stated in the following proposition. Throughout, let $p_{\star}$ denote a uniform lower bound on the dual-variable induced stationary distributions over the state-action space, [^3] An explicit expression for $p_{\star}$ is derived in the proof of Lemma 25. valid for all $\rho\in H$. Denote the minimal number of visits to any state-action pair until iteration k as $\underline{\nu}_{k}:=\min_{(s,a)\in\mathcal{X}}\nu_{k}(s,a)$. It holds that $\underline{\nu}_{k}\leq\nu_{k}(s,a)\leq k$ for all $(s,a)\in\mathcal{X}$. Denote the reduced Lagrangian objective as
@@
Recall that $\bar{\alpha}_{k}=\alpha(\tilde{\nu}_{k}(s_{k}))$ and $\bar{\beta}_{k}=\beta(\nu_{k}(X_{k}))$.
Proposition 15
@@
The proof of Proposition 15 is provided in Section C.1.
@@
$\tau_{k}^{2}\alpha_{\mathrm{env},k-\tau_{k}}\leq\tfrac{\tilde{\eta}_{V}}{18\,C_{1}\,C_{\mathrm{env}}},\hskip 2.0pt\tau_{k}\alpha_{\mathrm{env},k-\tau_{k}}\leq\min\{\tfrac{1}{\sqrt{3}D},\,\tfrac{1}{\sqrt{3}LB}\},\hskip 2.0pt\alpha_{k}\leq\tfrac{1}{\tilde{\eta}_{V}},\hskip 2.0pt\tfrac{p_{\star}k}{2}\geq 1,$ (24)
@@
Theorem 16 (Convergence rate of Algorithm 2)
Let Assumption 13 hold. Let $\{\rho_{k}\}$, $\{V_{k}\}$ denote the iterates of Algorithm 2 with the projection on the primal iterates to $\mathcal{V}_{r}$. Fix $\delta\in(0,1)$ and let $\mathcal{G}_{\delta}$, and $K(\delta)$ be as in Proposition 15. Let the stepsize schedules be as in (21) with
$\beta_{0}\geq\frac{2}{p_{\star}\mu_{\mathrm{opt}}},\ \ \frac{\beta_{0}}{\alpha_{0}}\leq\frac{\tilde{\eta}_{V}}{8p_{\star}\mu_{\mathrm{opt}}}.$
@@
$\mathbb{E}\!\left[f(\rho^{\star})-f(\rho_{k})|\mathcal{G}_{\delta}\right]$ $\leq\left(\mathbb{E}\!\left[f(\rho^{\star})-f(\rho_{k_{0}})|\mathcal{G}_{\delta}\right]+\frac{16L^{2}}{p_{\star}^{3}\tilde{\eta}_{V}}\frac{\beta_{k_{0}}}{\alpha_{k_{0}}}\mathbb{E}\!\left[\|V_{k_{0}}-\lambda(\rho_{k_{0}})\|_{2}^{2}|\mathcal{G}_{\delta}\right]\right)\frac{k_{0}+1}{k+1}$
$~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~+\frac{2C_{2}(\delta)\log^{2}(k+1)}{3(k+1)^{2/3}},$